OpenStudy (anonymous):

Assume that one number from 1 to 7 is equally likely to be selected at random. Each number corresponds to one of the seven figures shown. Three figures are (3) circles (2) triangles (2) square.Determine the probability of selecting a circle, given that a number greater than or equal to 5 is selected

hartnn (hartnn):

the probab of selecting '5' from 7 numbers is 1/7 'AND' the prob that '5' is on the circle is 3/7,so overall its,3/49....similarly for 6 and 7...so 3/49+3/49+3/49

OpenStudy (anonymous):

I'm not too sure what I did is right, but I'm actually doing this at school at the moment and I did not get the same answer as you hartnn. Can you tell me what I did wrong? Let Circle = A Let Number >/= 5 = B Pr(A) = 3/7 Pr(B) = 3/7 So I went P (A | B) = Pr (A and B) / Pr (B) Plugging in the numbers, I got the final answer to be 3/7. Hm.... :(

OpenStudy (fellowroot):

Is it not P(circle)=3/7 P(x>=5)=3/7 both need to be at the same time so its AND P(circle)P(x>=5)= 9/49

OpenStudy (fellowroot):

Nabsz, your formula is for a conditional probability. This means that one event has occurred and then another follows afterward. If the question had talked about getting the probability of the circle before you compare the number to <=5 then yes your formula works, but here I don't think order matters at all. So you just multiply both probabilities.

OpenStudy (anonymous):

But isn't the question saying that a number greater than or equal to 5 is selected and then a circle? I'm sorry for being a pain, I really don't understand probability very well. Thanks for explaining though Fellowroot :)

OpenStudy (fellowroot):

I think I see what you mean the word "then" seems like you need to do one first then the next one second, but it may not be in this case. A conditional probability would be stated for example: What is the probability that the first ball is yellow and next ball is yellow too. I'd like to see the correct answer just to check to see which answer is right.

OpenStudy (anonymous):

Yeah, I see what you mean too. This is why I dislike probability haha the answer sometimes depends on your interpretation of the question grr :(